# Self-Organization and Spectral Mechanism of Attractor Landscapes in High-Capacity Kernel Hopfield Networks

### v5

@@ 7 Conclusion
Our phenomenological analysis identified a “Ridge of Optimization” in the phase diagram, where the network achieves maximal stability characterized by a unique “Force Antagonism.” Crucially, our theoretical analysis revealed the spectral origin of this phenomenon: Spectral Concentration. We demonstrated that on the Ridge, the learning algorithm drives the weight spectrum into a critical state characterized by a massive leading eigenvalue for global stability, while retaining a non-zero heavy tail for high-capacity storage. This finding dispels the notion of a simple rank-1 collapse and instead portrays the optimal memory landscape as a finely tuned balance between attractive force and pattern discriminability.
@@ A.3 Theoretical Basis for Spectral Concentration (Proposition 2)
[…]
Akira TamamoriHe received his B.E., M.E., and D.E. degrees from Nagoya Institute of Technology, Nagoya, Japan, in 2008, 2010, 2014, respectively. From 2014 to 2016, he was a Research Assistant Professor at Institute of Statistical Mathematics, Tokyo, Japan. From 2016 to 2018, he was a Designated Assistant Professor at the Institute of Innovation for Future Society, Nagoya University, Japan. From 2018 to 2020, he was an lecturer at Aichi Institute of Technology, Japan. He has been an Associate Professor at Aichi Institute of Technology, Japan. He is a member of the Institute of Electronics, Information and Communication Engineers (IEICE), the Acoustical Society of Japan (ASJ), and Asia-Pacific Signal and Information Processing Association (APSIPA).

### v6

@@ 7 Conclusion
Our phenomenological analysis identified a Ridge of Optimization in the phase diagram, where the network achieves maximal stability characterized by a unique Force Antagonism. Crucially, our theoretical analysis revealed the spectral origin of this phenomenon: Spectral Concentration. We demonstrated that on the Ridge, the learning algorithm drives the weight spectrum into a critical state characterized by a massive leading eigenvalue for global stability, while retaining a non-zero heavy tail for high-capacity storage. This finding dispels the notion of a simple rank-1 collapse and instead portrays the optimal memory landscape as a finely tuned balance between attractive force and pattern discriminability.
@@
[…]
Not applicable.
[…]
The authors declare no competing interests.
[…]
All the authors contributed equally to the present work.
[…]
We used Gemini 2.5 Pro for proofreading the English manuscript.